OpenStudy (anonymous):

A doctor’s office schedules 15-minute appointments and half-hour appointments for weekdays. The doctor limits these appointments to, at most, 30 hours per week. Write an inequality to represent the number of 15-minute appointments x and the number of half-hour appointments y the doctor may have in a week.

OpenStudy (anonymous):

.25x + .5y < 30 .25 = 1/4 of an hour, or 15 minutes .5 = 1/2 of an hour, or 30 minutes